# Speeding up quantum measurement using space-time trade-off

### v1

@@ Abstract
We present a scheme for speeding up quantum measurement. The scheme builds on previous protocols that entangle the system to be measured with ancillary systems. In the idealised situation of perfect entangling operations and no decoherence, it gives an exact space-time trade-off meaning the readout time reduces linearly with the number of ancilla. We verify this scheme is robust against experimental imperfections through numerical modelling of gate noise and readout errors, and under certain circumstances our scheme can even lead to better than linear improvement in the speed of measurement with the number of systems measured. This hardware-agnostic approach is broadly applicable to a range of quantum technology platforms and offers a route to accelerate mid-circuit measurement as required for effective quantum error correction.
@@ II Case study: Poisson measurement statistics
In order to present the main idea, we will first focus on state dependent fluorescence qubit measurement which is the primary method used in trapped ion [2, 24] and neutral atom [1, 25, 26, 27] quantum computing.
Qubit information is encoded in either a “dark” or “bright” state of an ion or atom. When illuminated by a detection laser, the atom will be projected into one of these states and will then scatter photons. A measurement device (detector) will count the number $k$ of photons detected. We have in mind to count photons for a time $t$. The bright state will scatter many photons with a mean of $\nu_{1}(t)$ while the dark state will scatter relatively fewer photons with a mean of $\nu_{0}(t)$ where $\nu_{1}(t)\geq\nu_{0}(t)$. To discriminate the bright and dark states we define a threshold value $\eta(t)$ where $\nu_{0}(t)<\eta(t)<\nu_{1}(t)$. If the number of photons counted is $k<\eta(t)$ then we assign the outcome 00 and if $k\geq\eta(t)$ we assign the outcome $1$. From now on we refer to the bright state as |$\ket{1}$ and the dark state as |$\ket{0}$.
Let us denote by $\mathcal{P}_{\ket{j},t}^{(n)}(k)$ the probability distribution of the random variable $K_{\ket{j},t}^{(n)}$, defining to the number of photons counted, when we use $n$ detectors for time $t$ and the input state is |$\ket{j}$, $j\in\{0,1\}$. Initially we consider $n=1$.
@@
A key property of the Poisson distribution is that the parameter defining it is additive. If $k_{1}$ photons are emitted during time $t_{1}$ and $k_{2}$ photons are emitted during a subsequent time $t_{2}$ such that the total number of photons emitted in time $t=t_{1}+t_{2}$ is $k=k_{1}+k_{2}$, then by the fundamental convolution property of the Poisson distribution,
@@
Although one is generally interested in how good a measurement is for an arbitrary state, a commonly used figure of merit is the distinguishability, as quantified by the signal-to-noise-ratio (SNR), of the |$\ket{0}$ and |$\ket{1}$ state probability distributions. For two arbitrary probability distributions $P_{0}(x)$ and $P_{1}(x)$, the SNR is
$S(P_{0},P_{1})=\frac{2|\langle X\rangle_{P_{0}}-\langle X\rangle_{P_{1}}|}{\Delta X_{P_{0}}+\Delta X_{P_{1}}}.$ (2)
@@
In our case, for the two Poisson distributions in (1) we find
$S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})=\frac{2\sqrt{t}|\mu_{0}-\mu_{1}|}{\sqrt{\mu_{0}}+\sqrt{\mu_{1}}}.$ (3)
@@
II.1 Exact space-time trade-off in the noiseless case
@@
We now introduce a scheme which, rather than trading time for quality as above, allows us to trade time for “space” – by which we mean number of qubits. In particular, this provides a means to realise high-quality fast measurements, at the cost of using additional ancillas. Our scheme is outlined in Fig. 1, and employs a simple ‘cloning’ protocol. We apply a CNOT gate between our target qubit and an ancillary qubit initialised in the state |$\ket{0}$. We assume CNOTs can be performed in a negligibly-short time compared to the measurement duration. We measure the ancilla and the original qubit then assign the result as either 00 or $1$, representing the outcome of the measurement on the target qubit. In the ideal case, when measured for a time $t$, we will now show that this scheme gives identical measurement statistics to that of a standard measurement on a single qubit measured for a time $2t$. This thus effectively halves the measurement time.
@@
Figure 1: Space-time trade-off scheme. Example using one additional ancillary qubit. We entangle the input qubit with an ancillary qubit in state |$\ket{0}$ using a CNOT gate. We then measure each qubit for a time $t$, obtaining results $k_{1}$ and $k_{2}$ respectively. The statistics of the sum $k=k_{1}+k_{2}$ coincides with the statistics for a single measurement of time $2t$. We thus effectively speed up the measurement by a factor of 2 using this scheme.
@@
For now, for clarity of presentation, we analyse the case of independent detectors for each qubit, but come back to this point below. We denote by $k_{1}$ the number of photons detected by the 1st detector, and $k_{2}$ for the second. The measurement statistics for the pair of outcomes $(k_{1},k_{2})$ are
|$\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})=L_{\mu_{j}t}(k_{1})L_{\mu_{j}t}(k_{2}),$ (4)
@@
where |$\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})$ is the measurement statistics for the input state |$\ket{jj}$ when each measurement lasts for a time $t$. The outcomes $(k_{1},k_{2})$ form a pair of independent random variables, hence the probability distribution of their sum $k=k_{1}+k_{2}$ is the convolution of the individual probability distributions, hence
$\mathcal{P}_{\ket{j},t}^{(2)}(k)=(L_{\mu_{j}t}*L_{\mu_{j}t})(k)=L_{2\mu_{j}t}(k).$ (5)
@@
An important feature is present here; since the probability distribution only considers the total sum of emissions from both fluorescing qubits, we could instead consider the distribution as describing the number of photons counted from a single detector which is collecting all the emitted photons, rather than collecting separately and then adding.
The key insight is that, due to the convolution property of the Poisson, measuring a single qubit for time $2t$ is equivalent to summing the results from two measurements, each of time $t$, on two entangled qubits. Specifically
$\mathcal{P}_{\ket{j},t}^{(2)}(k)=\mathcal{P}_{\ket{j},2t}^{(1)}(k).$ (6)
@@
This result shows that we can speed the measurement up by a factor of $2$, by performing a statistically identical measurement in time $\frac{t}{2}$ as was previously performed in time $t$. Since the measurements are identical, the quality, as quantified by the SNR (or any other figure of merit), will be constant while the measurement is performed faster.
@@
where $k$ is the total sum of photons counted and $(f^{*N})(w)$ is the $N$-fold convolution of the function $f(v)$. This shows that we can keep the quality of a measurement constant, and speed it up arbitrarily, from time $t$ to time $t/N$, by making use of $N$ measurements in parallel.
A useful ratio for comparison with more realistic devices later on is $t^{(1)}_{S}/t^{(N)}_{S}$ where $t^{(N)}_{S}$ is the time it takes an $N$ qubit scheme to reach an SNR of $S$. In this case we have
@@
for all $S$, so the improvement in speed is exactly linear in qubit number. The speed up from our scheme is shown in Fig. 2 (a) and (b). The parameters used for the plots are experimentally relevant for neutral atom quantum computers and are found in [28].
III General space-time trade-off
A key insight is that the above analysis is much more general than just Poisson statistics arising from counting photons. In fact, the above scheme will lead to a space-time trade-off in a wide range of situations, with very few assumptions about the characteristics of the measuring device.
@@ III General space-time trade-off
For the case of Poisson statistics above, we saw that $\alpha~{}=~{}\frac{1}{2}$, however we may imagine an alternative dependence on time.
@@
Figure 2: Space-time trade-off for idealised and realistic devices. Plots show how faster measurements can be made without sacrificing distinguishability by introducing more qubits. The parameters used for the plots are experimentally relevant and are found in [28]. In plots a) and b), we use Poisson distributed measurement statistics (7) with $\mu_{0}=3.5\text{ms}^{-1}$ and $\mu_{1}=14\text{ms}^{-1}$ for different numbers of qubits $N$. a) shows the SNR (3) against time from $N=1$ to $N=5$ qubits. The SNR will always increase with time, however the rate at which the SNR improves increases with qubit number. The dashed grey lines are lines of constant SNR at $S=8,9.5,10.6$ and are plotted in b) which shows the time of measurement $t$ against number of qubits from $N=1$ to $N=10$ qubits. The inset shows the same points but for the inverse time $1/t$. The inset clearly shows the linear relationship implied by (8). In plots c) and d) we show similar results but for realistic devices. In this case, a higher maximum distinguishability can be achieved by introducing more qubits and faster measurements can be made while also improving distinguishability. These plots are generated using the complete noise model of the Poisson distributed measurement statistics (VIII.1), found in the Appendix, with the cascade compilation of noisy CNOT gates defined in the Supplementary material. We use input parameters $\mu_{0}=3.5\text{ms}^{-1}$, $\mu_{1}=14\text{ms}^{-1}$, $\lambda=0.0041\text{ms}^{-1}$ and $p=0.01$ where $p$ is chosen as a conservative estimate of CNOT failure rate. c) shows the SNR (2) against time for different numbers of qubits $N$. The SNR increases with time up to a maximum value then decreases due to decoherence. The rate at which the SNR improves increases with qubit number, as does the maximum SNR achieved. The dashed grey lines are lines of constant SNR at $S=8,9.5,10.6$ and are plotted in d) which shows the time of measurement $t$, when the defined SNR is first achieved, against number of qubits $N$. Notice how $S=9.5$ has no point for $N=1$ and $S=10.6$ has no points for $N=1,2$; this means these values cannot be reached without more qubits. The inset shows the same points plotted for inverse time $1/t$. The points do not follow a straight line so the space-time trade-off is not linear with qubit number $N$. The Appendix includes figures with extra details for c) and d). Fig. 3 demonstrates the increase in the peak SNR value with qubit number and Fig. 4 demonstrates the increase of the speed up with qubit number.
@@
Now, in the above, we saw that in an idealised setting, involving no noise in the scheme, that by bringing in $N-1$ additional ancilliary qubits, applying CNOTs and measuring each, that the statistics of the sum of outcomes is precisely the convolution of the statistics of an individual measurement, as given in (7). Here, while general distributions $\mathcal{P}_{\ket{j},t}^{(1)}$ do not convolve to distributions of the same form, we nevertheless have the fundamental property that the expectation and variance add,
$\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(1)}},$ $\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(1)}}.$ (10)
It therefore follows that the SNR for this measurement scheme will be
$S(\mathcal{P}_{\ket{0},t}^{(N)},\mathcal{P}_{\ket{1},t}^{(N)})=\sqrt{N}S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})\sim t^{\alpha}\sqrt{N}.$ (11)
@@
This recovers (8) in the case $\alpha=1/2$ but shows more generally, if $\alpha>1/2$, we will still achieve a trade-off between space and time using this scheme, albeit not as good as (8). However, if $\alpha<1/2$, we will achieve a “super-improvement” where the trade-off is better than (8).
One application of this general result is that it allows us to take into account decoherence in the measurement process. In the next section we address the important question of how imperfections in the scheme affect its overall performance.
IV Space-time trade-off for more realistic devices
@@
For our scheme to be useful in practical applications it is crucial to show that the speed-up is robust to realistic imperfection. To this end, we now describe a simple noise model to take into account these dominant sources of error. Since ancilla qubits are initialised in state |$\ket{0}$, imperfections in the CNOT gate operations can be modelled by considering errors on state |$\ket{10}$ being mapped to state |$\ket{00}$ with error probability p (see eg. [29]). Additionally, we consider imperfections in qubit readout caused by qubits in state |$\ket{1}$ falling into |$\ket{0}$ during the measurement, such as due to depuming by off-resonant scattering in atomic fluorescence measurements [30]. This error is included by applying a probabilistic decay at constant rate $\lambda$ during the measurement. See the Appendix for further details on our noise model.
The effects are summarised in Figures 2c and 2d, in direct comparison to the perfect case. The SNR is calculated using (29) in the Supplementary material, which gives the SNR of a realistic $N$ qubit scheme entirely in terms of the expectation and variance of the single qubit states |$\ket{0}$ and |$\ket{1}$ and the expectation and variance of the CNOT noise.
@@
We see that we still obtain a significant space-time trade-off, allowing for significantly faster measurements, even in the presence of noise, demonstrating the robustness of this scheme.
@@
There are other notable points worth highlighting. First, unlike in the perfect case, the SNR does not increase indefinitely. It reaches a peak before starting to decrease, due to decoherence. We see that the peak value is increased by schemes using more qubits. Thus not only does the scheme speed up the measurement while keeping the quality constant, it can also improve the measurement at the same time. This is shown in Fig. 3 in the Appendix in more detail. Similar to the perfect case, the rate of SNR increase improves with $N$, but it can now be a “super-improvement”; for small enough values of $N$ and CNOT noise $p$, the rate of speed up is better than (8) for the relevant time scales, that is the times before decoherence causes the measurement to deteriorate, as demonstrated by Fig. 4 in the Appendix.
V Conclusion

### v2

@@ Abstract
We present a scheme for speeding up quantum measurement. The scheme builds on previous protocols that entangle the system to be measured with ancillary systems. In the idealised situation of perfect entangling operations and no decoherence, it gives an exact space-time trade-off meaning the readout speed increases linearly with the number of ancilla. We verify this scheme is robust against experimental imperfections through numerical modelling of gate noise and readout errors, and under certain circumstances our scheme can even lead to better than linear improvement in the speed of measurement with the number of systems measured. This hardware-agnostic approach is broadly applicable to a range of quantum technology platforms and offers a route to accelerate mid-circuit measurement as required for effective quantum error correction.
@@ II Case study: Poisson measurement statistics
In order to present the main idea, we will first focus on state dependent fluorescence qubit measurement which is the primary method used in trapped-ion [2, 24] and neutral-atom [1, 25, 26, 27] quantum computing.
Qubit information is encoded in either a “dark” or “bright” state of an ion or atom. When illuminated by a detection laser, the atom will be projected into one of these states and will then scatter photons. A measurement device (detector) will count the number $k$ of photons detected. We have in mind to count photons for a time $t$. The bright state will scatter many photons with a mean of $\nu_{1}(t)$ while the dark state will scatter relatively fewer photons with a mean of $\nu_{0}(t)$ where $\nu_{1}(t)\geq\nu_{0}(t)$. To discriminate the bright and dark states we define a threshold value $\eta(t)$ where $\nu_{0}(t)<\eta(t)<\nu_{1}(t)$. If the number of photons counted is $k<\eta(t)$ then we assign the outcome 00 and if $k\geq\eta(t)$ we assign the outcome $1$. From now on we refer to the bright state as |$\ket{1}$ and the dark state as |$\ket{0}$. In all that follows, we denote $j\in\{0,1\}$.
Let us denote by $\mathcal{P}_{\ket{j},t}^{(n)}(k)$ the probability distribution of the random variable $K_{\ket{j},t}^{(n)}$, defining the number of photons counted, when we use $n$ detectors for time $t$ and the input state is |$\ket{j}$. Initially we consider $n=1$.
@@
If $k_{1}$ photons are emitted during time $t_{1}$ and $k_{2}$ photons are emitted during a subsequent time $t_{2}$ such that the total number of photons emitted in time $t=t_{1}+t_{2}$ is $k=k_{1}+k_{2}$, then by the fundamental convolution property of the Poisson distribution,
@@
Although one is generally interested in how good a measurement is for an arbitrary state, a commonly used figure of merit is the distinguishability, as quantified by the signal-to-noise-ratio (SNR), of the ${\mathcal{P}_{\ket{0},t}^{(n)}(k)}$ and ${\mathcal{P}_{\ket{1},t}^{(n)}(k)}$ probability distributions. The SNR is
$S\left(\mathcal{P}_{\ket{0},t}^{(n)},\mathcal{P}_{\ket{1},t}^{(n)}\right)=\frac{2\left|\langle X\rangle_{\mathcal{P}_{\ket{0},t}^{(n)}}-\langle X\rangle_{\mathcal{P}_{\ket{1},t}^{(n)}}\right|}{\Delta X_{\mathcal{P}_{\ket{0},t}^{(n)}}+\Delta X_{\mathcal{P}_{\ket{1},t}^{(n)}}}.$ (2)
@@
In our case, for the two Poisson distributions, we find
$S\left(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)}\right)=\frac{2\sqrt{t}|\mu_{0}-\mu_{1}|}{\sqrt{\mu_{0}}+\sqrt{\mu_{1}}}.$ (3)
@@
III Exact space-time trade-off in the idealised Poisson case
@@
We now introduce a scheme which, rather than trading time for quality as above, allows us to trade time for “space” – by which we mean number of qubits. In particular, this provides a means to realise high-quality fast measurements, at the cost of using additional ancillas. See Fig. 1. We apply a CNOT gate between our target qubit and an ancillary qubit initialised in the state |$\ket{0}$. In this section we consider the idealised situation that the CNOTs can be performed perfectly and in a negligibly-short time compared to the measurement duration. We measure the ancilla and the original qubit then assign the result as either 00 or $1$, representing the outcome of the measurement on the target qubit. In the ideal case, when measured for a time $t$, we will now show that this scheme gives identical measurement statistics to that of a standard measurement on a single qubit measured for a time $2t$. The measurement time is effectively halved.
@@
Figure 1: Space-time trade-off scheme. Idealised version using one additional ancillary qubit. We entangle the input qubit with an ancillary qubit in state |$\ket{0}$ using a CNOT gate. We then measure each qubit for a time $t$, obtaining results $k_{1}$ and $k_{2}$ respectively. The statistics of the sum $k=k_{1}+k_{2}$ coincides with the statistics for a single measurement of time $2t$. We thus speed up the measurement by a factor of 2.
@@
For now, for clarity of presentation, we analyse the case of independent detectors for each qubit, but come back to this point below. We denote by $k_{1}$ and $k_{2}$ the number of photons detected from the first and second qubit detectors respectively and we input the state |$\ket{\psi}=\ket{j}$ into our scheme. The measurement statistics for the pair of outcomes $(k_{1},k_{2})$, given we are measuring the post-CNOT state |$\ket{jj}$ for a time $t$, are
|$\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})=L_{\mu_{j}t}(k_{1})L_{\mu_{j}t}(k_{2}).$ (4)
@@
The outcomes $(k_{1},k_{2})$ form a pair of independent random variables, hence the probability distribution of their sum $k=k_{1}+k_{2}$ is the convolution of the individual probability distributions, thus ${\mathcal{P}_{\ket{j},t}^{(2)}(k)=(L_{\mu_{j}t}*L_{\mu_{j}t})(k)=L_{2\mu_{j}t}(k)}$.
@@
An important feature is present here; since the probability distribution only considers the total sum of emissions from both fluorescing qubits, we can consider the distribution as describing the number of photons counted from a single detector which is collecting all the emitted photons, rather than collecting separately and then adding.
@@
The key insight is that, due to the convolution property of the Poisson, measuring a single qubit for time $2t$ is equivalent to summing the results from two measurements, of time $t$, on two entangled qubits. Specifically ${\mathcal{P}_{\ket{j},t}^{(2)}(k)=\mathcal{P}_{\ket{j},2t}^{(1)}(k).}$ This result shows that we can speed up the measurement, by a factor of $2$, by performing a statistically identical measurement in time $\frac{t}{2}$ as was previously performed in time $t$. Since the measurements are identical, the quality, as quantified by the SNR (or any other figure of merit), will be constant while the measurement is performed faster.
@@
where $k$ is the total number of photons counted and $(f^{*N})(w)$ is the $N$-fold convolution of the distribution $f(v)$. This shows that we can keep the quality of a measurement constant, and speed it up arbitrarily, from time $t$ to time $t/N$, by making use of $N$ measurements in parallel.
A useful ratio for comparison with more realistic devices later on is $t^{(1)}_{S}/t^{(N)}_{S}$ where $t^{(N)}_{S}$ is the time it takes an $N$ qubit scheme to reach an SNR of $S$. In this case
@@
for all $S$, so the improvement in speed is exactly linear in qubit number. The speed up from our scheme is shown in Figs. 2a and 2b.
IV General space-time trade-off
A key insight is that the scheme above applies much more generally than just to Poisson statistics arising from counting photons. In fact, the scheme gives a space-time trade-off in a wide range of measuring devices.
@@ IV General space-time trade-off
For the case of Poisson statistics above we saw that $\alpha~{}=~{}1/2$ however we may imagine an alternative dependence on time.
@@
Figure 2: Space-time trade-off for idealised and realistic devices. The parameters used for the plots are experimentally relevant and are taken from [28]. In plots a) and b), we use Poisson distributed measurement statistics (5) with $\mu_{0}=3.5\text{ms}^{-1}$ and $\mu_{1}=14\text{ms}^{-1}$ for different numbers of qubits $N$. a) SNR (2) against time from $N=1$ to $N=5$ qubits. The SNR always increases with time and qubit number and continues the trend beyond what is shown on the graph. b) time of measurement $t$ against number of qubits from $N=1$ to $N=10$ qubits for the SNR values on the grey dashed lines in a). Inset: same points but for the inverse time $1/t$, demonstrating the linear relationship (6). c) and d) show similar results but for realistic devices. In this case, a higher maximum distinguishability can be achieved by introducing more qubits and faster measurements can be made while also improving distinguishability. These plots are generated using the noise and decoherence model (IX), found in the Appendix. We use input parameters $\lambda=0.0041\text{ms}^{-1}$ and $p=0.01$ where $p$ is chosen as a conservative estimate of CNOT failure rate. In c), the SNR increases with time up to a maximum value then decreases due to decoherence. The rate at which the SNR improves increases with qubit number, as does the peak SNR achieved. Inset: points in the inset no longer follow a straight line, indicating that the space-time trade-off is not linear with qubit number $N$.
@@
In the above we saw that, in an idealised version of the scheme involving perfect CNOTs, by bringing in $N-1$ additional ancilliary qubits, applying CNOTs and measuring each, that the statistics of the sum of outcomes is precisely the convolution of the statistics of an individual measurement, as given in (5). Here, while general distributions $\mathcal{P}_{\ket{j},t}^{(1)}$ do not convolve to distributions of the same form, for perfect CNOTs we nevertheless have the fundamental property that the expectation and (due to independence) variance add,
$\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(1)}},$ $\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(1)}}.$
It therefore follows that the SNR, using the scheme, will be
$S\left(\mathcal{P}_{\ket{0},t}^{(N)},\mathcal{P}_{\ket{1},t}^{(N)}\right)=\sqrt{N}S\left(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)}\right)\sim t^{\alpha}\sqrt{N}.$ (8)
@@
This recovers (6) in the case $\alpha=1/2$ but shows more generally, if $\alpha>1/2$, we will still achieve a trade-off between space and time using this scheme, albeit not as good as (6). Nevertheless, if $\alpha<1/2$, we will achieve a “super-improvement” where the trade-off is better than (6).
V Space-time trade-off in more realistic settings
@@
For our scheme to be useful in practical applications it is crucial to show that the speed-up is robust to realistic imperfection. To this end, we now describe a simple noise model to take into account the dominant sources of error. Since ancilla qubits are initialised in state |$\ket{0}$, imperfections in the CNOT gate operations can be modelled by considering errors on state |$\ket{10}$ being mapped to state |$\ket{00}$ with error probability p (see eg. [29]). Additionally, we consider imperfections in qubit readout caused by qubits in state |$\ket{1}$ falling into |$\ket{0}$ during the measurement, such as due to depuming by off-resonant scattering in atomic fluorescence measurements [30]. This error is included by applying a probabilistic decay at constant rate $\lambda$ during the measurement. See the Appendix for further details on our noise model.
@@
The effects are summarised in Figs. 2c and 2d. The SNR is calculated using (26) in the Supplementary Material, which gives the SNR of a realistic $N$ qubit scheme entirely in terms of the expectation and variance of the number of photons counted for an $N=1$ qubit scheme $K_{\ket{j},t}^{(1)}$ and the expectation and variance of the CNOT noise probability distribution. We see that we still obtain a significant space-time trade-off, allowing for significantly faster measurements, even in the presence of noise, demonstrating the robustness of this scheme.
We note that our discussions have been framed in terms of the SNR, but we find essentially the same space-time trade-off if we consider the measurement infidelity: this is discussed in the Supplementary Material. In the perfect case this is shown analytically while for the imperfect case we present numerical results analogous to Fig. 2c and 2d.
@@
There are other notable points worth highlighting. First, unlike in the perfect case, the SNR does not increase indefinitely. It reaches a peak before starting to decrease, due to decoherence. We see that the peak value is increased by schemes using more qubits. Thus not only does the scheme speed up the measurement while keeping the quality constant, it can also improve the measurement at the same time. This is shown in Fig. 3 in the Appendix in more detail. Similar to the perfect case, the rate of SNR increase improves with $N$, but it can now be a “super-improvement”; for small enough values of $N$ and CNOT failure rate $p$, the rate of speed up is better than linear (6) for the relevant time scales, that is the times before decoherence causes the measurement to deteriorate, as demonstrated by Fig. 4 in the Appendix.
VI Conclusion